OpenStudy (anonymous):

my problem is 3/4w-1/2w-4=12?

OpenStudy (anonymous):

SOLVE FOR w ?????

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

FIRST ADD 4 to both sides

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

i got 64

OpenStudy (anonymous):

3/4w-1/2w-4=12 3/4w-1/2w-4+4=12+4 (3/4-1/2)w=16 {(3-2)/4}w=16 w/4 =16 w=64

OpenStudy (anonymous):

U r \[\checkmark\]